OpenStudy (anonymous):

What is 1+2+3+4+5+6+7+8+9+10.

OpenStudy (anonymous):

Just answer for an easy medal.

OpenStudy (anonymous):

I'll give you a hint, it is 55.

OpenStudy (anonymous):

Ooh that's a tough one. Have you tried adding them together? See if 55 works out for you :)

OpenStudy (anonymous):

You got the medal, good job.

OpenStudy (anonymous):

Lol Thanks :)

OpenStudy (anonymous):

I was trying to see if someone would try for the easy medal.

OpenStudy (anonymous):

It took two hours.

OpenStudy (anonymous):

sum of (1 to n) = (n+1)(n)/2 = (11)(10)/2 = 55